OpenStudy (paige):

A curve is such that dy/dx= 2Cos(2x-π/2). The curve passes through the point (π/2,3). Find the equation of the curve & the equation of the normal to the curve at the pt. where x=3π/4.

OpenStudy (anonymous):

To get the equation of the curve you need to integrate that function with respect to x. \[\int\limits_{}^{}2\cos(2x-\frac{\pi}{2})dx= sin(2x-\frac{\pi}{2})+c\] Knowing that the point (pi/2, 3) is on the curve, and substituting those values in the equation we can figure out what c is: \[3= \sin(2(\frac{\pi}{2})-\frac{\pi}{2})+c \Rightarrow 3 = \sin(\frac{\pi}{2})+c\Rightarrow 3 = 1+c\] So c is 2, thus the equation of the curve must be: \[f(x) = \sin(2x-\frac{\pi}{2})+ 2\]

OpenStudy (anonymous):

You can integrate it easily. Then you have: \[\sin(2x-\frac{\pi}{2})+C\] Using the point you can solve for C. You can plug in your x value into dy/dx, get your slope. Thats the slope of the tangent. That tells you that the slope of the normal is perpendicular or the negative reciprocal of it. From there, you have a slope and a point, you can use point slope form to get a line. PSF: \[y-y_0=m(x-x_0)\] :)

OpenStudy (paige):

Thank you. Can you tell me the equation of the normal to the curve at the pt. where x=3π/4 ?

OpenStudy (anonymous):

Well, plugging in 3pi/4 you get that m=-2. Taking the negative reciprocal you have m=1/2. So you have y-3=(1/2)(x-pi/2) y-3=(1/2)x-pi/4 y=(1/2)x-pi/4-3

OpenStudy (paige):

Where do you 'plug' in 3π/4 ?

OpenStudy (anonymous):

Into x for dy/dx So you have: \[\frac{dy}{dx}|_{x=\frac{3\pi}{4}}=2\cos(2(\frac{3\pi}{4})-\frac{\pi}{2})=2\cos(\pi)=2(-1)=-2\]

OpenStudy (paige):

Got the answer!! Thank you so much! :)

OpenStudy (anonymous):

No problem :)